A 10.0 L gas sample is heated to 400 ◦C at P = constant. The final volume is 20.0 L. What is the initial temperature of this gas? Assume ideal-gas behavior and a constant amount of gas throughout the process (n, P = constant). T(K) = T(◦C) + 273.

(A) 63.5 ◦C (B) 73 ◦C (C) 200 ◦C (D) 336.5 ◦C (E) 609.5 ◦C
Answer #1

Given:

Tf = 400 oC

= (400+273) K

= 673 K

Vi = 10.0 L

Vf = 20.0 L

use:

Vi/Ti = Vf/Tf

10.0 L / Ti = 20.0 L / 673.0 K

Ti = 336.5 K

= (336.5 - 273) oC

= 63.5 oC

Answer: A